OpenStudy (anonymous):

6^-2 6^-2 @Mathematics

OpenStudy (anonymous):

6^-2 = 1/6^2 = 1/36

OpenStudy (anonymous):

Thanks, is -2^4=+/-16?

OpenStudy (anonymous):

No. -2^4 is (-2)(-2)(-2)(-2), which equals 16.

OpenStudy (anonymous):

So is x^4 +9=90 x=3?

OpenStudy (anonymous):

Yes x = 3

OpenStudy (anonymous):

x^4+9=90, x^4=90-9, x^4=81, x=81^1/4, x=3